OpenStudy (anonymous):

explanation would be awesome:) In a right-angled triangle, the two shorter side lengths are 3 cm and 4 cm. To the nearest degree, the value of the smallest angle is

OpenStudy (anonymous):

Recall that the hypotenuse is always going to be the LONGEST side in a right triangle. Therefore, the shorter sides are 3 cm and 4 cm. Now, there are different ways to do this and this is only one: We can use the Tan ratio to determine one of the angles. Tanx = 3/4 x = 36.9 = 37 Since a triangle's angles add up to 180 degrees... 37 + 90 + x = 180 Using algebra... 180 - 37 - 90 = 53 degrees. Now that we found all angles, we find out that the smallest angle is 37. Therefore, 37. It would be best to find ALL angles before deciding which would be larger, I think. Just for the save of showing all work.

OpenStudy (anonymous):

So the answer would be....what @sky202

OpenStudy (anonymous):

53 degrees

OpenStudy (anonymous):

no answer would be 37

OpenStudy (anonymous):

the smalest angle would be 37

OpenStudy (anonymous):

oh yeah got it now thanks so much :)

OpenStudy (anonymous):

you are ever so welcome

OpenStudy (anonymous):

:)